Ali is a student taking statistics course unfortunately Ali is not a good student Ali does not read the book before class and does not do homework and misses classes. Ali intend to rely on quiz. Quiz consist of 10 multiple choice questions.each have 5 possible answer and only one is correct.
(a) What is the probability that Ali gets no correct answer?
(b) What is the probability that Ali gets 2 correct answe?
(c) What is the probability that Ali gets all correct answer?
(d) What is the probability that Ali gets 2 to 5 correct answer?

Answer by Boreal(15235)

p(correct)=0.2, p(incorrect)=0.8
all 10 incorrect is .8^10=0.1074
all correct is .2^10=1.024 x 10 ^(-7)
2 correct is 10C2*0.2^2*0.8^8=0.3020
2 to 5 correct
for 3 it is 10C3*0.2^3*0.8^7=0.2013
for 4 it is 10C4*0.2^4*0.8^6=0.0881
for 5 it is 10C5*0.2^5*0.8^5=0.0264
total is 0.6178